# Detection of evolutionary shifts in variance under an Ornstein–Uhlenbeck model

### v1

@@ 7 Data accessibility
The tree of Anolis lizards is provided by Mahler et al. (2013). The phylogenetic tree and principal component data of cordylid data ( Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016) are available in the R package phytools: http://www.phytools.org/Rbook/. Our R package is available at https://github.com/WenshaZ/ShiVa.
Acknowledgement
@@
Appendices
A Supplemental figures
Figure 18: The false positive rates of different methods for detecting shifts in optimal value when there are only shifts in variance (Shift: 197; 198; 192)
Figure 19: The predictive log-likelihood difference of different methods when there are only shifts in variance (Shift: 197; 198; 192)

### v2

@@ Appendix B Data accessibility
The tree of Anolis lizards is provided by Mahler et al. (2013b). It can be found in the Dryad data repository: https://doi.org/10.5061/dryad.9g182. The phylogenetic tree and principal component data of cordylid data (Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016b) can be found in the Dryad data repository: https://doi.org/10.5061/dryad.k186f. Our R package is available at https://github.com/WenshaZ/ShiVa.
Appendix C Author contributions
All authors contributed to conceive the ideas, design the methodology and the simulations. WZ conducted the simulations, analyzed the results and wrote the manuscript under the guidance of LSTH and TK. LSTH and TK reviewed and edited the manuscript.
Appendix D Acknowledgement